# Stochastic Particle Acceleration during Pressure-Anisotropy-Driven Magnetogenesis in the Pre-Structure Universe

### 05 Mar 2026 v1

@@ II.2 Implications for Pre-Shock CR Production
$c_{s}(z)=\sqrt{\frac{\gamma k_{B}T(z)}{\mu m_{p}}}$ (11)
is the adiabatic sound speed of the background plasma. We adopt $\mu\approx 0.59$ appropriate for a fully ionized primordial plasma. Observations and simulations of weakly magnetized astrophysical plasmas, such as the intracluster medium, suggest that gas motions are typically subsonic relative to the sound speed. For example, the Hitomi satellite measured a velocity dispersion of $\sim 160\,{\rm km\,s^{-1}}$ in the Perseus cluster core, indicating subsonic gas motions [13], while numerical studies likewise find turbulent velocities that are a fraction of the sound speed [8]. Motivated by these results, we explore a representative range $\epsilon=0.05$$0.5$.
Pressure anisotropy and plasma beta.
@@
$\Delta(z)\approx\Delta_{0}=\text{const}.$ (12)
@@
$\beta(z)\sim\frac{2}{|\Delta_{0}|},$ (13)
@@
$\Delta_{\rm ex}(\delta,\beta)\equiv|\Delta_{0}+\delta\Delta|-|\Delta_{0}|=|\Delta_{0}+\delta\Delta|-2\beta^{-1},$ (14)
which reduces to $\Delta_{\rm ex}=\delta\Delta$ for $\Delta_{0}>0$ and $\delta\Delta\geq 0$. In high-$\beta$ astrophysical plasmas such as the intracluster medium (ICM), where $\beta\sim\mathcal{O}(10^{2})$ [e.g., 30, 31, 32, 27, 17, 18], this corresponds to a typical anisotropy level $\Delta_{0}\sim\mathcal{O}(10^{-2})$.
Magnetic-field evolution.
@@
$\frac{dB}{dt}$ $\approx\omega_{\rm cp,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}$
$+\omega_{\rm cp,0}\frac{|\delta\Delta|^{3/2}}{1+|\Delta|}e^{-\Gamma_{d}t-\frac{1}{2}\Gamma_{d}/\Gamma_{c}^{(\rm inst)}}\left(\frac{B}{B_{0}}\right)^{2},$ (15)
where the second term represents the contribution from perturbations of the pressure anisotropy. Here $\Gamma_{d}$ denotes the effective damping rate of the anisotropy due to pitch-angle scattering by plasma instabilities, whereas $\Gamma_{c}^{(\rm inst)}$ represents the characteristic rate at which large-scale plasma motions generate pressure anisotropy. $B_{0}$ is the seed magnetic field and $\omega_{\rm cp,0}$ corresponds to the proton gyrofrequency evaluated at $B_{0}$.
@@
$\frac{dB}{dt}\approx\omega_{\rm cp,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}.$ (16)
@@
II.3 Analytic constraint on the critical magnetic field
A necessary condition for efficient cosmological acceleration is that the acceleration time be shorter than the Hubble time, $t_{\rm acc}<t_{H}=H^{-1}(z)$. Combining Eqs.  (4), (9), and (14) yields an analytic lower bound on the magnetic field:
$B>B_{\rm crit}(z)\equiv\frac{m_{p}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (17)
Eq. (17) makes explicit the strong parametric dependence $B_{\rm crit}\propto H(z)\,\epsilon^{-2}c_{s}^{-2}(z)\Delta_{\rm ex}^{-3/2}$: unless the turbulent motions are sufficiently strong and/or the anisotropy is strongly supercritical, the magnetic field required for $t_{\rm acc}<t_{H}$ becomes unrealistically large, implying that pressure-anisotropy-driven instabilities alone are unlikely to produce a significant pre-shock CR population under generic cosmological conditions.
Using Eq. (17), we can derive an explicit redshift dependence of the critical field required for $t_{\rm acc}<t_{H}$. At $z\gtrsim 1$ (matter-dominated era) we approximate
$H(z)\simeq H_{0}\sqrt{\Omega_{m}}\,(1+z)^{3/2}.$ (18)
@@ II.3 Analytic constraint on the critical magnetic field
$T(z)=T_{0}\left(\frac{1+z}{1+z_{0}}\right)^{\alpha},$ (19)
the sound speed satisfies $c_{s}^{2}(z)\propto T(z)\propto(1+z)^{\alpha}$. Eq. (17) then yields
$B_{\rm crit}(z)\propto\frac{H(z)}{c_{s}^{2}(z)}\propto(1+z)^{\frac{3}{2}-\alpha},$ (20)
@@
$B_{\rm crit}(z)=B_{\rm crit}(z_{0})\,\left(\frac{1+z}{1+z_{0}}\right)^{\frac{3}{2}-\alpha},$ (21)
where $B_{\rm crit}(z_{0})$ collects the normalization factors, including $\epsilon^{-2}$ and $\Delta_{\rm ex}^{-3/2}$. For the illustrative case $\alpha=1$ used in our toy thermal model, $B_{\rm crit}(z)\propto(1+z)^{1/2}$, implying a slowly increasing critical field toward higher redshift.
Figure 1: Redshift evolution of the magnetic field $B(z)$ (black) compared with the analytic critical field $B_{\rm crit}(z)$ (red). Their intersection defines the CR turn-on redshift $z_{\rm on}$. The curves are shown for fiducial parameters $B_{0}=10^{-18}$ G, $\delta\Delta/\Delta_{0}=0.1$, $\Delta_{0}\sim 2\beta^{-1}=10^{-2}$, and $\epsilon=0.2$.
Figure 2: Dependence of the CR turn-on redshift $z_{\rm on}$ on (a) the turbulence amplitude $\epsilon$ and (b) the pressure-anisotropy perturbation $\delta\Delta$. All other parameters are fixed to the fiducial values used in Fig. 1. Over the explored parameter range, $z_{\rm on}$ varies only weakly, remaining close to $z_{\rm on}\sim 1.7$.
II.4 Definition of the CR turn-on redshift
A useful way to quantify when pressure-anisotropy-driven scattering can begin to enable cosmological stochastic acceleration is to define a “CR turn-on redshift” $z_{\rm on}$. We define $z_{\rm on}$ as the epoch at which the stochastic acceleration time becomes comparable to the Hubble time:
$\left.\frac{t_{\rm acc}(z)}{t_{H}(z)}\right|_{z=z_{\rm on}}=1,\qquad t_{H}(z)=H^{-1}(z).$ (22)
For $z>z_{\rm on}$ the expansion time is too short ($t_{\rm acc}>t_{H}$) and stochastic acceleration is inefficient, while for $z<z_{\rm on}$ acceleration becomes viable ($t_{\rm acc}<t_{H}$).
@@ II.4 Definition of the CR turn-on redshift
$B(z_{\rm on})=B_{\rm crit}(z_{\rm on}),$ (23)
where $B_{\rm crit}(z)$ is defined by Eq. (17). In the matter-dominated regime $H(z)\propto(1+z)^{3/2}$ and for a parametrized thermal history $T(z)\propto(1+z)^{\alpha}$ (so that $c_{s}^{2}\propto(1+z)^{\alpha}$), one finds $B_{\rm crit}(z)\propto(1+z)^{3/2-\alpha}$.
For illustrative purposes, if the magnetic-field evolution can be approximated by $B(z)=B_{0}(1+z)^{p}$, the turn-on redshift follows from Eq. (23) as
$1+z_{\rm on}=\left(\frac{C}{B_{0}}\right)^{\!\frac{1}{\,p-\left(\frac{3}{2}-\alpha\right)}},$ (24)
@@
To illustrate this behavior explicitly, we compute the magnetic-field evolution $B(z)$ using the unperturbed scattering-dominated solution from our previous work and compare it with the analytic critical field $B_{\rm crit}(z)$. Fig. 1 shows the resulting redshift evolution for a representative set of parameters. At high redshift the magnetic field remains far below the critical value, implying $t_{\rm acc}\gg t_{H}$ and therefore inefficient stochastic acceleration. As cosmic time progresses, nonlinear magnetic amplification sets in and $B(z)$ increases rapidly. The CR turn-on redshift $z_{\rm on}$ is defined by the crossing $B(z)=B_{\rm crit}(z)$. In the fiducial case shown here, this occurs at $z_{\rm on}\sim 1.7$, corresponding to a magnetic field strength of order $10^{-9}$ G. Importantly, this epoch coincides with the onset of nonlinear structure formation, suggesting that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven stochastic processes.
@@
Fig. 2 illustrates the dependence of the turn-on redshift on the turbulence amplitude $\epsilon$ and on the pressure anisotropy, parameterized by $\delta\Delta/\Delta_{0}$ and $\Delta_{0}\sim 2\beta^{-1}$. Despite the explicit appearance of both quantities in the analytic expression for $B_{\rm crit}(z)$, the resulting variation in $z_{\rm on}$ is remarkably small over the considered parameter range. In particular, even when $\epsilon$ and $\delta\Delta$ are varied by factors of a few, the turn-on redshift remains confined to a narrow interval around $z_{\rm on}\sim 1.7$. This behavior reflects the fact that $B(z)$ exhibits a rapid nonlinear growth phase, so that modest vertical shifts in $B_{\rm crit}(z)$ translate into only small horizontal shifts in the crossing redshift. The onset of cosmologically viable stochastic acceleration is therefore robust against moderate changes in the underlying microphysical parameters. Importantly, the inferred $z_{\rm on}$ remains close to the epoch of nonlinear structure formation, reinforcing the conclusion that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven processes.
Figure 3: Redshift evolution of the maximum attainable proton energy $E_{\max}(z)$ computed from Eq. (28) for turbulence amplitudes $\epsilon=0.05$ (black), $0.1$ (red), and $0.2$ (blue). Near the turn-on epoch $z_{\rm on}\sim 1.7$, $E_{\max}$ reaches $\sim 10^{2}\,\mathrm{GeV}$.
II.5 Maximum attainable energy during cosmological magnetogenesis
@@
While the condition $t_{\rm acc}(z)<t_{H}(z)$ determines when stochastic acceleration becomes cosmologically viable, it does not specify the maximum particle energy that can be reached. To quantify the physical significance of the turn-on epoch, we estimate the maximum proton energy attainable before the expansion of the Universe limits further acceleration.
Acceleration time: optimistic upper bound.
@@
To estimate the largest particle energies that could in principle be reached, we adopt an optimistic upper bound for the particle mean free path by assuming $\lambda\sim r_{L}$, where $r_{L}$ is the Larmor radius. This corresponds to the strong-scattering limit $\nu_{\rm scatt}\sim\omega_{\rm cp}$, which can arise when instability-driven pitch-angle scattering dominates. In this limit, the spatial diffusion coefficient becomes
$D(E)\sim\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (25)
@@
for relativistic protons with $E\approx pc$. In this limit, the second-order Fermi acceleration time can be expressed as
$t_{\rm acc,2}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{r_{L}(E)}{c}=\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{E}{eB(z)c}.$ (26)
@@
Definition of the maximum energy.
@@
$t_{\rm acc,2}(E_{\max},z)=t_{H}(z)=H^{-1}(z).$ (27)
Substituting the parametrization $v_{\rm tur}(z)=\epsilon c_{s}(z)$ into Eq. (26), the maximum attainable energy becomes
$E_{\max}(z)\approx eB(z)c\epsilon^{2}\left(\frac{c_{s}(z)}{c}\right)^{2}H^{-1}(z).$ (28)

### 01 Apr 2026 v2

@@ 2.2 Implications for Pre-Shock CR Production
$c_{s}(z)=\sqrt{\frac{\gamma k_{B}T(z)}{\mu m_{p}}}$ (2.14)
is the adiabatic sound speed of the background plasma. We adopt $\mu\approx 0.59$ appropriate for a fully ionized primordial plasma. Observations and simulations of diffuse astrophysical plasmas, such as the intracluster medium, suggest that gas motions are typically subsonic relative to the sound speed. For example, the Hitomi satellite measured a velocity dispersion of $\sim 160,{\rm km,s^{-1}}$ in the Perseus cluster core, indicating subsonic turbulence [14], while numerical studies likewise find turbulent velocities that are a fraction of the sound speed [8]. Although these results pertain to gas that has been processed by structure-formation shocks in the low-redshift Universe, they provide a useful reference for the characteristic magnitude of turbulent motions in weakly magnetized plasmas. Motivated by this, we adopt a representative range $\epsilon=0.05$$0.5$ in the present study.
@@
$\Delta(z)\approx\Delta_{0}=\text{const}.$ (2.15)
@@
$\beta(z)\sim\frac{2}{|\Delta_{0}|},$ (2.16)
@@
$\Delta_{\rm ex}(\delta,\beta)\equiv|\Delta_{0}+\delta\Delta|-|\Delta_{0}|=|\Delta_{0}+\delta\Delta|-2\beta^{-1},$ (2.17)
which reduces to $\Delta_{\rm ex}=\delta\Delta$ for $\Delta_{0}>0$ and $\delta\Delta\geq 0$. In high-$\beta$ astrophysical plasmas such as the intracluster medium (ICM), where $\beta\sim\mathcal{O}(10^{2})$ [e.g., 32, 33, 34, 28, 18, 19], this corresponds to a typical anisotropy level $\Delta_{0}\sim\mathcal{O}(10^{-2})$.
@@
$\frac{dB}{dt}$ $\approx\omega_{\rm ci,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}$
$+\omega_{\rm ci,0}\frac{|\delta\Delta|^{3/2}}{1+|\Delta|}e^{-\Gamma_{d}t-\frac{1}{2}\Gamma_{d}/\Gamma_{c}^{(\rm inst)}}\left(\frac{B}{B_{0}}\right)^{2},$ (2.18)
where the second term represents the contribution from perturbations of the pressure anisotropy. Here $\Gamma_{d}$ denotes the effective damping rate of the anisotropy due to pitch-angle scattering by plasma instabilities, whereas $\Gamma_{c}^{(\rm inst)}$ represents the characteristic rate at which large-scale plasma motions generate pressure anisotropy. $B_{0}$ is the seed magnetic field and $\omega_{\rm ci,0}$ corresponds to the ion gyrofrequency evaluated at $B_{0}$.
@@
$\frac{dB}{dt}\approx\omega_{\rm ci,0}\frac{\left(|\Delta|-2\beta^{-1}\right)^{3/2}}{1+|\Delta|}\left(\frac{B}{B_{0}}\right)^{2}.$ (2.19)
@@
2.3 Analytic constraint on the critical magnetic field
A necessary condition for efficient cosmological acceleration is that the acceleration time be shorter than the Hubble time, $t_{\rm acc}<t_{H}=H^{-1}(z)$. Combining Eqs.  (2.5), (2.10), and (2.17) yields an analytic lower bound on the magnetic field:
$B>B_{\rm crit}(z)\approx\frac{m_{i}c}{e}\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\left(\frac{v(E)}{c}\right)\frac{H(z)}{\Delta_{\rm ex}^{3/2}}.$ (2.20)
This condition defines a critical magnetic field required for stochastic acceleration to compete with cosmological expansion, and therefore represents a necessary (but not sufficient) criterion for particle acceleration. Eq. (2.20) makes explicit the strong parametric dependence $B_{\rm crit}\propto H(z)\,\epsilon^{-2}c_{s}^{-2}(z)\Delta_{\rm ex}^{-3/2}$: unless the turbulent motions are sufficiently strong and/or the anisotropy is strongly supercritical, the magnetic field required for $t_{\rm acc}<t_{H}$ becomes unrealistically large, implying that pressure-anisotropy-driven instabilities alone are unlikely to produce a significant pre-shock CR population under generic cosmological conditions.
Using Eq. (2.20), we can derive an explicit redshift dependence of the critical field required for $t_{\rm acc}<t_{H}$. At $z\gtrsim 1$ (matter-dominated era) we approximate
$H(z)\simeq H_{0}\sqrt{\Omega_{m}}\,(1+z)^{3/2}.$ (2.21)
@@ 2.3 Analytic constraint on the critical magnetic field
$T(z)=T_{0}\left(\frac{1+z}{1+z_{0}}\right)^{\alpha},$ (2.22)
the sound speed satisfies $c_{s}^{2}(z)\propto T(z)\propto(1+z)^{\alpha}$. Eq. (2.20) then yields
$B_{\rm crit}(z)\propto\frac{H(z)}{c_{s}^{2}(z)}\propto(1+z)^{\frac{3}{2}-\alpha},$ (2.23)
@@
$B_{\rm crit}(z)=B_{\rm crit}(z_{0})\left(\frac{1+z}{1+z_{0}}\right)^{\frac{3}{2}-\alpha},$ (2.24)
where $B_{\rm crit}(z_{0})$ collects the normalization factors, including $\epsilon^{-2}$ and $\Delta_{\rm ex}^{-3/2}$. The parameter $\alpha$ characterizes the thermal history of the background plasma through $T(z)\propto(1+z)^{\alpha}$. In the absence of heating processes, adiabatic expansion of non-relativistic particles leads to $T\propto a^{-2}$, corresponding to $\alpha=2$. However, in realistic cosmological environments, additional heating mechanisms—such as photoheating, residual Compton heating, and structure formation—can partially offset adiabatic cooling, resulting in a shallower temperature evolution with $\alpha<2$. For this reason, it is reasonable to consider a range of thermal histories with $1\lesssim\alpha\lesssim 2$. In the illustrative case $\alpha=1$, corresponding to a moderately heated medium, the critical field scales as $B_{\rm crit}(z)\propto(1+z)^{1/2}$, implying a slowly increasing threshold toward higher redshift. By contrast, for purely adiabatic evolution ($\alpha=2$), one obtains $B_{\rm crit}(z)\propto(1+z)^{-1/2}$.
2.4 Definition of the CR turn-on redshift
A useful way to quantify when pressure-anisotropy-driven scattering can begin to enable cosmological stochastic acceleration is to define a ”CR turn-on redshift” $z_{\rm on}$. We define $z_{\rm on}$ as the epoch at which the stochastic acceleration time becomes comparable to the Hubble time,
$\left.\frac{t_{\rm acc}(z)}{t_{H}(z)}\right|_{z=z_{\rm on}}=1,\qquad t_{H}(z)=H^{-1}(z).$ (2.25)
This condition provides a necessary criterion for stochastic acceleration on cosmological timescales. For $z>z_{\rm on}$ the expansion time is too short ($t_{\rm acc}>t_{H}$) and stochastic acceleration is inefficient, while for $z<z_{\rm on}$ acceleration may in principle become viable. However, this condition is not sufficient, as additional energy loss processes—such as Coulomb and ionization losses—can impose more stringent constraints, particularly at low energies. Therefore, the $z_{\rm on}$ defined here should be interpreted as an upper bound on the redshift below which stochastic acceleration may operate.
Figure 1: Redshift evolution of the magnetic field $B(z)$ (black) compared with the analytic critical field $B_{\rm crit}(z)$ for two different thermal histories: $\alpha=1$ (solid red) and $\alpha=2$ (dashed red), corresponding to a moderately heated medium and purely adiabatic evolution, respectively. Their intersection defines the CR turn-on redshift $z_{\rm on}$. The curves are shown for fiducial parameters $B_{0}=10^{-18}$ G, $\delta\Delta/\Delta_{0}=0.1$, $\Delta_{0}\sim 2\beta^{-1}=10^{-2}$, and $\epsilon=0.2$. Here, the critical field is evaluated in the relativistic limit ($v\sim c$), providing a reference estimate for particle acceleration. The resulting $z_{\rm on}$ depends only weakly on the assumed thermal history.
Figure 2: Dependence of the CR turn-on redshift $z_{\rm on}$ on (a) the turbulence amplitude $\epsilon$ and (b) the pressure-anisotropy perturbation $\delta\Delta$. All other parameters are fixed to the fiducial values used in Figure 1. Over the explored parameter range, $z_{\rm on}$ varies only weakly, remaining close to $z_{\rm on}\sim 1.7$.
@@ 2.4 Definition of the CR turn-on redshift
$B(z_{\rm on})=B_{\rm crit}(z_{\rm on}),$ (2.26)
where $B_{\rm crit}(z)$ is defined by Eq. (2.20). In the matter-dominated regime $H(z)\propto(1+z)^{3/2}$ and for a parametrized thermal history $T(z)\propto(1+z)^{\alpha}$ (so that $c_{s}^{2}\propto(1+z)^{\alpha}$), one finds $B_{\rm crit}(z)\propto(1+z)^{3/2-\alpha}$. For illustrative purposes, if the magnetic-field evolution can be approximated by $B(z)=B_{0}(1+z)^{p}$, the turn-on redshift follows from Eq. (2.26) as
$1+z_{\rm on}=\left(\frac{C}{B_{0}}\right)^{\!\frac{1}{\,p-\left(\frac{3}{2}-\alpha\right)}},$ (2.27)
@@
To illustrate this behavior explicitly, we compute the magnetic-field evolution $B(z)$ using the unperturbed scattering-dominated solution from our previous work and compare it with the analytic critical field $B_{\rm crit}(z)$. Here, the critical field is evaluated in the relativistic limit ($v\sim c$), providing a baseline estimate for particle acceleration. Figure 1 shows the resulting redshift evolution for a representative set of parameters, including two different thermal histories characterized by $\alpha=1$ and $\alpha=2$. At high redshift, the magnetic field remains far below the critical value in both cases, implying $t_{\rm acc}\gg t_{H}$ and therefore inefficient stochastic acceleration. As cosmic time progresses, nonlinear magnetic amplification driven by pressure-anisotropy instabilities leads to a rapid growth of $B(z)$. The CR turn-on redshift $z_{\rm on}$ is defined by the crossing $B(z)=B_{\rm crit}(z)$, marking the epoch at which stochastic acceleration becomes viable on cosmological timescales. In all cases considered here, the crossing occurs at $z_{\rm on}\sim 1.7$, corresponding to a magnetic field strength of order $10^{-9}$ G. Notably, the location of this crossing depends only weakly on the assumed thermal history. This can be understood from the different redshift scalings of the two quantities: while $B_{\rm crit}(z)$ evolves only as a power law in $(1+z)$, the magnetic field undergoes rapid nonlinear amplification. As a result, the onset of stochastic acceleration is primarily governed by instability-driven magnetization, with the thermal history playing only a secondary role. Importantly, this epoch coincides with the onset of nonlinear structure formation, suggesting that efficient cosmic-ray production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven stochastic processes.
@@
Figure 2 illustrates the dependence of the turn-on redshift on the turbulence amplitude $\epsilon$ and on the pressure anisotropy, parameterized by $\delta\Delta/\Delta_{0}$ and $\Delta_{0}\sim 2\beta^{-1}$. Despite the explicit appearance of both quantities in the analytic expression for $B_{\rm crit}(z)$, the resulting variation in $z_{\rm on}$ is remarkably small over the considered parameter range. In particular, even when $\epsilon$ and $\delta\Delta$ are varied by factors of a few, the turn-on redshift remains confined to a narrow interval around $z_{\rm on}\sim 1.7$. This behavior reflects the fact that $B(z)$ exhibits a rapid nonlinear growth phase, so that modest vertical shifts in $B_{\rm crit}(z)$ translate into only small horizontal shifts in the crossing redshift. The onset of cosmologically viable stochastic acceleration is therefore robust against moderate changes in the underlying microphysical parameters. Importantly, the inferred $z_{\rm on}$ remains close to the epoch of nonlinear structure formation, reinforcing the conclusion that efficient CR production is naturally associated with structure-formation shocks rather than with earlier microinstability-driven processes.
2.5 Maximum attainable energy during cosmological magnetogenesis
@@
While the condition $t_{\rm acc}(z)<t_{H}(z)$ determines when stochastic acceleration becomes cosmologically viable, it does not specify the maximum particle energy that can be reached. To quantify the physical significance of the turn-on epoch, we estimate the maximum ion energy attainable before the expansion of the Universe limits further acceleration.
@@
To estimate the largest particle energies that could in principle be reached, we adopt an optimistic upper bound for the particle mean free path by assuming $\lambda\sim r_{L}$, where $r_{L}$ is the Larmor radius. This corresponds to the strong-scattering limit $\nu_{\rm scatt}\sim\omega_{\rm ci}$, which can arise when instability-driven pitch-angle scattering dominates. In this limit, the spatial diffusion coefficient becomes
$D(E)\sim\frac{1}{3}r_{L}(E)c,\qquad r_{L}(E)=\frac{pc}{eB}\approx\frac{E}{eB},$ (2.28)
@@
for relativistic ions with $E\approx pc$. In this limit, the second-order Fermi acceleration time can be expressed as
$t_{\rm acc,2}(E,z)\approx\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{r_{L}(E)}{c}=\left(\frac{c}{v_{\rm tur}(z)}\right)^{2}\frac{E}{eB(z)c}.$ (2.29)
@@
Figure 3: Redshift evolution of the maximum attainable ion energy $E_{\max}(z)$ computed from Eq. (2.31) for turbulence amplitudes $\epsilon=0.05$ (black), $0.1$ (red), and $0.2$ (blue). Near the turn-on epoch $z_{\rm on}\sim 1.7$, $E_{\max}$ reaches $\sim 10^{2}\,\mathrm{GeV}$.
@@
$t_{\rm acc,2}(E_{\max},z)=t_{H}(z)=H^{-1}(z).$ (2.30)
Substituting the parametrization $v_{\rm tur}(z)=\epsilon c_{s}(z)$ into Eq. (2.29), the maximum attainable energy becomes
$E_{\max}(z)\approx eB(z)c\epsilon^{2}\left(\frac{c_{s}(z)}{c}\right)^{2}H^{-1}(z).$ (2.31)